# finite free probability and $S$ transforms of Jacobi processes

### v1

@@ Abstract.
In this paper, we study the $S$ transforms of Jacobi processes in the frameworks of free and finite free probability theories. We begin by deriving a partial differential equation satisfied by the free $S$ transform of the free Jacobi process, and we provide a detailed analysis of its characteristic curves. We turn next our attention to the averaged characteristic polynomial of the Hermitian Jacobi process and to the dynamic of its roots, referred to as the frozen Jacobi process. In particular, we prove, for a specific set of parameters, that the former aligns up to a Szegö variable transformation with the Hermite unitary polynomial. We also provide an expansion of the averaged characteristic polynomial of the Hermitian process in the basis of Jacobi polynomials. Finally, we establish the convergence of the frozen Jacobi process to the free Jacobi process in high dimensions by using the finite free S transform. In doing so, we prove a general result, interesting in its own, on the convergence of the finite differences of the finite free $S$ transform.
@@ 1.1. Finite free probability (FFP)
Introduced and studied in [MSS22], FFP has gained a lot of attention in recent years, see [AP18, Ari+24, MSS22, ALR25]. It provides an original perspective on certain operations on polynomials that were defined long before the seminal work [MSS22]. Actually, FFP involves convolution operations on averaged characteristic polynomials of complex random matrices, whose laws have prescribed symmetry invariances, as well as the computation of their root distributions. Important questions deal with asymptotics of these root distributions when the degrees of their corresponding polynomials tend to infinity. In this regime, the FFP convolutions tend to the convolutions already introduced by Voiculescu in his theory of free probability. Further investigations concern finite free counterparts of results and transforms of prime importance in Voiculescu’s theory.
In [MSS22], finite versions of the free additive, the free rectangular and the free multiplicative convolutions were studied, though the latter is attributed to [Mir21]. The finite free unitary convolution was also defined in [Mir21], and the corresponding central limit theorem was proved there. In this respect, the limiting polynomial is given by the so-called unitary Hermite polynomial, which was shown in [Kab25] to coincide (up to rescaling) with the characteristic polynomial of the Brownian motion on the unitary group. In that paper, it was also proved that the corresponding empirical root distribution converges weakly to the spectral distribution of the free unitary Brownian motion [Bia97].
@@ 1.2. Crystallization and the $\beta$-Jacobi process
Among finite versions of various results in random matrix and free probability theories figure the analytic extrapolation to any $\beta>0$ of finite free convolutions and the $\beta$-corners processes [GM20]. In particular, this extrapolation shows that the elementary symmetric polynomials in the roots of averaged characteristic polynomials of random matrices with prescribed law-invariance properties do not depend on $\beta$. This fact was already noticed in earlier papers for classical Hermite/Laguerre/Jacobi ensembles and opens the way to take the limit $\beta\rightarrow+\infty$. Doing so leads to the extension of the so-called crystallization beyond the eigenvalues of classical matrix-valued processes.
@@
The $\beta$-Jacobi process, introduced and studied in [Dem10], is an extrapolation to all $\beta>0$ of the eigenvalues of the real symmetric and the complex Hermitian Jacobi processes. For any $\beta>0$, its distribution converges weakly as time $t\rightarrow+\infty$ to the corresponding $\beta$-Jacobi ensemble. Besides, when $\beta=2$, the underlying Hermitian matrix model converges in the large-size limit to the free Jacobi process [Dem08]. The moments of this matrix model and their large size limits were computed in a series of papers [DD18, DHS20, DH25]. However, the Lebesgue decomposition of its spectral distribution was determined in [Ham18]. Additionally, the expressions for the $R$ and $S$ transforms of the free Jacobi process associated with a specific parameter set were derived in [Dem17] using the Lagrange inversion formula. In this respect, note that partial differential equations for $R$ and $S$ transforms relative to hydrodynamic limits of Dyson and Wishart processes were derived in [EKS22] and solved using the method of characteristics.
@@
As to the crystallization of the $\beta$-Jacobi process, it still holds as shown in [Voi23]. Yet, it does not follow from the approach undertaken in [GM20] since the marginal distribution of this random particle system does not satisfy the required invariance properties. Rather, a combination of various results proved in [Voi20], [Voi23] and [Voi25] shows that the averaged elementary symmetric polynomials in the rescaled process at any fixed time $t$ do not depend on $\beta$. Equivalently, the averaged characteristic polynomial coincides with that of the deterministic Jacobi particle system, referred to as the frozen Jacobi process. From a dynamical perspective, the differential system satisfied by the latter is equivalent to the inverse Jacobi-heat equation for the former [Voi25]. Apparently, such equivalence was noticed for the first time by T. Tao [^1] https://terrytao.wordpress.com/2017/10/17/heat-flow-and-zeroes-of-polynomials/ for polynomial roots undergoing the heat flow and was motivated by an application to analytic number theory. A few years later, other root dynamics were derived in the unpublished paper [Mar22], which relied on determinantal formulas describing various finite free convolutions. Actually, these dynamics are governed by the drifts of eigenvalue processes of the Dyson, Laguerre and Hermitian Jacobi processes. However, the underlying matrix model from which the last dynamics arise is not the Hermitian Jacobi process as explained in [Mar22] (see section 5.2). For the sake of completeness, let us close this paragraph by recalling the following related results:
@@
  • The long-time behaviour of the frozen Dyson, Laguerre and Jacobi processes is governed by the zeroes of Hermite, Laguerre and Jacobi polynomials (see e.g. [Voi25]).
@@
  • The dynamics of the roots of the Hermite unitary polynomial (which lie on the unit circle) were determined in [Mir21]. In particular, their angles form the frozen process corresponding to the unitary $\beta$-Dyson process.
1.3. finite free S transforms
@@
To clarify connections between finite free probability and Voiculescu’s theory, finite equivalents of various transforms central to free probability have been defined, including the $S$-transform. The first proposal for the finite free $S$ transform is due to Marcus [MSS22], followed a few years later by another one [Ari+24], motivated by a TCL for finite free multiplicative convolution. Marcus’s definition is general enough to cover both unitary and positive distributions, but is much less practical for computations. For this reason, we prefer to work with the definition proposed in [Ari+24] in the present work.
@@ 1.4. Our contributions
In the present work, we prove out-of-equilibrium equivalents of the previously stated results. More precisely, we investigate the $S$-transforms of Jacobi processes within the frameworks of both free and finite free probability theories. We begin by deriving a partial differential equation (PDE) that the $S$-transform of the free Jacobi process satisfies, as stated in Theorem 3.1.1. Once we do, we provide a detailed analysis of its stationary analytic solutions around the origin in relation to the angle operator between two free orthogonal projections, viewed as an operator in the compressed probability space. For the sake of simplicity and motivated by [Dem17], we determine the characteristic curves of this PDE for the special parameter values $\lambda=1$ and $\theta=1/2$ and obtain an explicit expression of the S transform locally around the origin.
Our focus is then turned to the averaged characteristic polynomial of the Hermitian Jacobi process and to the dynamics of its roots. In this respect, we write a direct proof of the fact that the former satisfies the inverse Jacobi-heat equation and prove further that the latter are contained in the interval $(0,1)$. In particular, when $\lambda=1$ and $\theta=1/2$, we show that the averaged characteristic polynomial at time $t$ aligns, up to a Szegö type transformation, with the Hermite unitary polynomial at time $2t$. This result is indeed the finite free analogue of Corollary 3.3 from [DHH12]. Additionally, we provide an expansion of the averaged characteristic polynomial of the Hermitian process in the basis of Jacobi polynomials, valid for any initial data. To this end, we appeal to the dual Cauchy identity for multivariate Jacobi polynomials and to the semi-group density of the Jacobi eigenvalue process. Doing so allows us to use the mutual orthogonality of these polynomials, leading to the sought expansion.
Finally, we establish that the crystallization dynamics of the Jacobi process tend, in the regime of high degree and high dimension, towards the free Jacobi process and write two proofs of this result. The first one relies on the equi-continuity of the moment sequence of the empirical root distribution and on the fact that the sequence of the moments of the free Jacobi process is uniquely determined by the differential system it satisfies. While the second proof is based on the finite free equivalent of the $T$-transform in Voiculescu’s theory. This alternative approach is noteworthy because it requires a general technical result that we prove below regarding the convergence of the finite $T$-transform and its discrete derivative. Once proved, we derive a differential-difference equation for the $T$-transform of the Hermitian Jacobi process and recover in the large size regime, the previously derived PDE for the free S transform.
@@ 2.1. The Hermitian Jacobi process
Let $(Y_{t}(d))_{t\geq 0}$ be a unitary Brownian motion of size $d$ and variance $t$. Let $P_{m},Q_{p}$ be two diagonal orthogonal projections of ranks $m\leq p$. Then the Hermitian Jacobi process $(X_{t}(m))_{t\geq 0}$ of parameters $(r,s,m)$ is defined by:
@@
where $(B_{t}^{i},t\geq 0)_{1\leq i\leq m}$ are independent Brownian motions [DD18]. Up to the time change $t\mapsto 2t$, this random particle system is an instance of the $\beta$-Jacobi process studied in [Dem08] corresponding to $\beta=2$. In particular, Corollary 2.1 shows that the eigenvalues process is globally defined in time provided that $p\wedge q>m-1/2$. Corollary 7 in [GM14] shows that the eigenvalues of $X_{t}$ are almost surely distinct and that global existence holds even under the weaker condition $p\wedge q>m-1$. Consequently, Theorem 4.1 in [Dem08] extends to $p\wedge q>m-1$ so that the probability distribution of the eigenvalues process has a density with respect to Lebesgue measure in $\mathbb{R}^{m}$.
@@ 2.2. The free Jacobi process
The free Jacobi process was introduced in [Dem08] as the large-size limit of the Hermitian Jacobi process $(X_{t/d(m)}(m))_{t\geq 0}$ in the asymptotic regime:
@@
where $P,Q$ are free (in Voiculescu’s sense) orthogonal projections in a non commutative probability space $(\mathcal{A},\tau)$ with ranks $\tau(P)=\lambda\theta$ and $\tau(Q)=\theta$, and $(Y_{t})_{t\geq 0}$ is the so-called free unitary Brownian motion [Bia97]. Viewed as an operator in the compressed probability space $(P\mathcal{A}P,\tau/\tau(P))$, the Cauchy transform $G_{t}^{(\lambda,\theta)}$ of its spectral distribution satisfies the following non-linear partial differential equation (PDE):
@@
The solution of this equation is subject to the initial data
@@
was determined in [DHH12] and used in [Dem17] to derive expansions of the corresponding $R$ and $S$ transforms, based on Lagrange inversion formula.
Other degenerate limiting regimes were investigated in [AVW24] for the Jacobi eigenvalue process, and gave rise to Marchenko-Pastur and Wigner distributions. At the root level, these regimes correspond to the ones studied in [DS95].
@@ 3.2. Stationary analytic (around the origin) solutions of (4)
This is the $S$ transform $S_{Q}$ of an orthogonal projection of rank $\theta$ taken at $\lambda\theta z$. Equivalently, it is the $S$ transform of the weak limit as $t\rightarrow+\infty$ of the spectral distribution of the free Jacobi process. The latter may be defined as the spectral distribution of the angle operator $PQP$ between two free orthogonal projections $P$ and $Q$ in a non commutative probability space $(\mathscr{A},\tau)$ with ranks $\tau(P)=\lambda\theta$ and $\tau(Q)=\theta$, viewed as an operator in the compressed probability space $(P\mathscr{A}P,\tau/\tau(P))$. Using the Nica-Speicher convolution semi-group, see [NS06], Lecture 14, the Voiculescu’s $R$ transform $\mathcal{R}_{PQP}$ of $PQP$ in the compressed space is given by
@@
where $R_{Q}$ is the Voiculescu’s $R$ transform of $Q$ in $(\mathscr{A},\tau)$, (see [NS06], page 211). Let
@@
be the shifted Voiculescu R transforms (also known as free cumulant generating functions) of $PQP$ and of $Q$, respectively. Then ([NS06], Lecture 16)

### v3

@@ Abstract.
In this paper, we study the $S$ transforms of Jacobi processes in the frameworks of free and finite free probability theories. We begin by deriving a partial differential equation satisfied by the free $S$ transform of the free Jacobi process, and we provide a detailed analysis of its characteristic curves. We turn next our attention to the averaged characteristic polynomial of the Hermitian Jacobi process and to the dynamic of its roots, referred to as the frozen Jacobi process. In particular, we prove, for a specific set of parameters, that the former aligns up to a Szegö variable transformation with the Hermite unitary polynomial. We also provide an expansion of the averaged characteristic polynomial of the Hermitian process in the basis of Jacobi polynomials. Finally, we establish the convergence of the frozen Jacobi process to the free Jacobi process in high dimensions by using the finite free S transform. In doing so, we prove a general result, interesting in its own, on the convergence of the finite differences of the finite free $S$ transform, which paves the way to obtain asymptotics of differential-difference equations satisfied by time-dependent finite free S transforms of polynomial sequences with positive roots.
@@ 1.1. Finite free probability (FFP)
Introduced and studied in [MR4408504], FFP has gained a lot of attention in recent years, see [arizmendi2018cumulants, arizmendi2024s, MR4408504, MR4925017]. It provides an original perspective on certain operations on polynomials that were defined long before the seminal work [MR4408504]. Actually, FFP involves convolution operations on averaged characteristic polynomials of complex random matrices, whose laws have prescribed symmetry invariances, as well as the computations of their root distributions. Important questions deal with asymptotics of these root distributions when the degrees of their corresponding polynomials tend to infinity. In this regime, the FFP convolutions tend to the convolutions already introduced by Voiculescu in his theory of free probability. Further investigations concern finite free counterparts of results and transforms of prime importance in Voiculescu’s theory (see e.g. [MR4925017], [MR4912666], [arizmendi2024s], [JKM25], [Kab25]).
In [MR4408504], finite versions of the free additive, the free rectangular and the free multiplicative convolutions were studied, though the latter is attributed to [MR4285332]. The finite free unitary convolution was also defined in [MR4285332], and the corresponding central limit theorem was proved there. In this respect, the limiting polynomial is given by the so-called unitary Hermite polynomial, which was shown in [MR4912666] to coincide (up to rescaling) with the characteristic polynomial of the Brownian motion on the unitary group. In that paper, it was also proved that the corresponding empirical root distribution converges weakly to the spectral distribution of the free unitary Brownian motion [MR1430721].
@@ 1.2. Crystallization and the $\beta$-Jacobi process
Among finite versions of various results in random matrix and free probability theories figure the analytic extrapolation to any $\beta>0$ of finite free convolutions and the $\beta$-corners processes [MR4073197]. In particular, this extrapolation shows that the elementary symmetric polynomials in the roots of averaged characteristic polynomials of random matrices with prescribed law-invariance properties do not depend on $\beta$. This fact was already noted in earlier papers for classical Hermite/Laguerre/Jacobi ensembles and opens the way to taking the limit $\beta\rightarrow+\infty$. Doing so leads to the extension of the so-called crystallization beyond the eigenvalues of classical matrix-valued processes.
@@
The $\beta$-Jacobi process, introduced and studied in [MR2719370], is an extrapolation to all $\beta>0$ of the eigenvalues of the real symmetric and the complex Hermitian Jacobi processes. For any $\beta>0$, its distribution converges weakly as time $t\rightarrow+\infty$ to the corresponding $\beta$-Jacobi ensemble. Besides, when $\beta=2$, the underlying Hermitian matrix model converges in the large-size limit to the free Jacobi process [MR2384475]. The moments of this matrix model and their large size limits were computed in a series of papers [MR3842169, MR4173022, MR4954166]. However, the Lebesgue decomposition of its spectral distribution was determined in [MR3812867]. Additionally, the expressions for the $R$ and $S$ transforms of the free Jacobi process associated with a specific parameter set were derived in [MR3679612] using the Lagrange inversion formula. In this respect, note that partial differential equations for $R$ and $S$ transforms relative to hydrodynamic limits of Dyson and Wishart processes were derived in [MR4446587] and solved using the method of characteristics.
@@
As to the crystallization of the $\beta$-Jacobi process, it still holds as shown in [MR4673381]. Yet, it does not follow from the approach undertaken in [MR4073197] since the marginal distribution of this random particle system does not satisfy the required invariance properties. Rather, a combination of various results proved in [MR4092353], [MR4673381] and [MR4781073] shows that the averaged elementary symmetric polynomials in the rescaled process at any fixed time $t$ do not depend on $\beta$. Equivalently, the averaged characteristic polynomial coincides with that of the deterministic Jacobi particle system, referred to as the frozen Jacobi process. From a dynamical perspective, the differential system satisfied by the latter is equivalent to the inverse Jacobi-heat equation for the former [MR4781073]. Apparently, such equivalence was noticed for the first time by T. Tao [^1] https://terrytao.wordpress.com/2017/10/17/heat-flow-and-zeroes-of-polynomials/ for polynomial roots undergoing the heat flow and was motivated by an application to analytic number theory. A few years later, other root dynamics were derived in the unpublished paper [FFPP], which relied on determinantal formulas describing various finite free convolutions. Actually, these dynamics are governed by the drifts of eigenvalue processes of the Dyson, Laguerre and Hermitian Jacobi processes. However, the underlying matrix model from which the last dynamics arise is not the Hermitian Jacobi process as explained in [FFPP] (see section 5.2). For the sake of completeness, let us close this paragraph by recalling the following related results:
@@
  • The long-time behaviour of the frozen Dyson, Laguerre and Jacobi processes is governed by the zeroes of Hermite, Laguerre and Jacobi polynomials (see e.g. [MR4781073]).
@@
  • The dynamics of the roots of the Hermite unitary polynomial (which lie on the unit circle) were determined in [MR4285332]. In particular, their angles form the frozen process corresponding to the unitary $\beta$-Dyson process.
1.3. (Finite) free S transforms
@@
To clarify connections between finite free probability and Voiculescu’s theory, finite analogues of various transforms central to free probability have been defined, including the $S$-transform [bercovici1993free]. The first proposal for the finite free $S$- transform is due to Marcus [MR4408504], followed a few years later by another proposal, motivated by the CLT for finite free multiplicative convolution [arizmendi2024s]. Marcus’s definition is general enough to cover the finite free multiplicative convolutions of distributions supported on the positive half-line and on the circle, but is much less practical for computations. For this reason, we prefer to work with the definition proposed in [arizmendi2024s] in the present work.
Recall that if $\mu$ is a probability measure on $\mathbb{R}_{+}$ different from the dirac measure at 0 and $\psi_{\mu}\colon\mathbb{C}_{+}\to\mathbb{C}$ is the moment transform of $\mu$ defined over the complex plane minus the positive half-line, then $\psi_{\mu}$ is univalent on the left upper-half plane $i\mathbb{C}_{+}$. Besides $\psi_{\mu}(i\mathbb{C}_{+})$ is a domain contained in the disc with diameter $(-1+\mu({0}),0)$ and $(-1+\mu({0}),0)\subset\psi_{\mu}(i\mathbb{C}_{+})$. The free $S$-transform of $\mu$ by, for $z$ in the vicinity of $(-1+\mu({0}),0)$ is then defined by
$S_{\mu}(z)=\frac{1+z}{z}\psi_{\mu}^{(-1)}(z).$
Moreover, it factorizes under the free multiplicative convolution of distributions on the positive half-line.
@@ 1.4. Our contributions
In the present work, we prove out-of-equilibrium equivalents of the previously stated results. More precisely, we investigate the $S$-transforms of Jacobi processes within the frameworks of both free and finite free probability theories. We begin by deriving a partial differential equation (PDE) that the $S$-transform of the free Jacobi process satisfies, as stated in Theorem 3.1.1. Once we do, we provide a detailed analysis of its stationary analytic solutions around the origin in relation to the angle operator between two free orthogonal projections, viewed as an operator in the compressed probability space. For the sake of simplicity and motivated by [MR3679612], we determine the characteristic curves of this PDE for the special parameter values $\lambda=1$ and $\theta=1/2$ and obtain an explicit expression of the S transform locally around the origin.
Our focus is then turned to the averaged characteristic polynomial of the Hermitian Jacobi process and to the dynamics of its roots. In this respect, we write a direct proof of the fact that the former satisfies the inverse Jacobi-heat equation and prove further that the latter are contained in the interval $(0,1)$. In particular, when $\lambda=1$ and $\theta=1/2$, we show that the averaged characteristic polynomial at time $t$ aligns, up to a Szegö type transformation, with the Hermite unitary polynomial at time $2t$. This result is indeed the finite free analogue of Corollary 3.3 from [MR3071702]. Additionally, we provide an expansion of the averaged characteristic polynomial of the Hermitian process in the basis of Jacobi polynomials, valid for any initial data. To this end, we appeal to the dual Cauchy identity for multivariate Jacobi polynomials and to the semi-group density of the Jacobi eigenvalue process. Doing so allows us to use the mutual orthogonality of these polynomials, leading to the sought expansion.
Finally, we establish that the crystallization dynamics of the Jacobi process tend, in the regime of high degree and high dimension, towards the free Jacobi process and write two proofs of this result. The first one relies on the equi-continuity of the moment sequence of the empirical root distribution and on the fact that the sequence of the moments of the free Jacobi process is uniquely determined by the differential system it satisfies. While the second proof is based on the finite free equivalent of the $T$-transform in Voiculescu’s theory. This alternative approach is noteworthy because it requires a general technical result that we prove below regarding the convergence of the finite $T$-transform and of its discrete derivative. Once proved, we derive a differential-difference equation for the $T$-transform of the Hermitian Jacobi process and recover in the large size regime, the previously derived PDE for the free S transform.
@@ 2.1. The Hermitian Jacobi process
Let $(Y_{t}(d))_{t\geq 0}$ be a unitary Brownian motion of size $d$ and variance $t$. Let $P_{m}$ and $Q_{p}$ be two diagonal orthogonal projections of rank $m$ and $p$ respectively, with $m\leq p$. The Hermitian Jacobi process $(X_{t}(m))_{t\geq 0}$ of parameters $(r,s,m)$ is the $m\times m$ matrix-valued stochastic process defined by:
@@
where $(B_{t}^{i},t\geq 0)_{1\leq i\leq m}$ are independent Brownian motions [MR3842169]. Up to the time change $t\mapsto 2t$, this random particle system is an instance of the $\beta$-Jacobi process studied in [MR2384475] corresponding to the value $\beta=2$. In particular, Corollary 2.1 shows that the eigenvalues process is globally defined in time provided that $p\wedge q>m-1/2$. Corollary 7 in [MR3296535] shows that the eigenvalues of $X_{t}$ are almost surely distinct and that global existence holds even under the weaker condition $p\wedge q>m-1$. Consequently, Theorem 4.1 in [MR2384475] extends to $p\wedge q>m-1$ so that the probability distribution of the eigenvalues process $(\lambda_{t}^{i},t\geq 0)_{1\leq i\leq m}$ has a density with respect to Lebesgue measure in $\mathbb{R}^{m}$, for any $p\wedge q>m-1$.
@@ 2.2. The free Jacobi process
The free Jacobi process was introduced in [MR2384475] as the large-size limit of the Hermitian Jacobi process $(X_{t/d(m)}(m))_{t\geq 0}$ in the asymptotic regime:
@@
where $P,Q$ are free (in Voiculescu’s sense) orthogonal projections in a non commutative probability space $(\mathcal{A},\tau)$ with ranks $\tau(P)=\lambda\theta$ and $\tau(Q)=\theta$, and $(Y_{t})_{t\geq 0}$ is the so-called free unitary Brownian motion [MR1430721]. Viewed as an operator in the compressed probability space $(P\mathcal{A}P,\tau/\tau(P))$, the Cauchy transform $G_{t}^{(\lambda,\theta)}$ of its spectral distribution satisfies the following non-linear partial differential equation (PDE):
@@
The solution of this equation subject to the initial data
@@
was determined in [MR3071702] and used in [MR3679612] to derive expansions of the corresponding $R$ and $S$ transforms, based on Lagrange inversion formula.
Other degenerate limiting regimes were investigated in [MR4751304] for the Jacobi eigenvalue process, and gave rise to Marchenko-Pastur and Wigner distributions. At the root level, these regimes correspond to the ones studied in [MR1342385].
@@ 3.2. Stationary analytic (around the origin) solutions of (4)
This is the $S$ transform $S_{Q}$ of an orthogonal projection $Q$ of rank $\theta$ taken at $\lambda\theta z$. Equivalently, it is the $S$ transform of the weak limit as $t\rightarrow+\infty$ of the spectral distribution of the free Jacobi process. The latter may be defined as the spectral distribution of the angle operator $PQP$ between two free orthogonal projections $P$ and $Q$ in a non commutative probability space $(\mathscr{A},\tau)$ with ranks $\tau(P)=\lambda\theta$ and $\tau(Q)=\theta$, viewed as an operator in the compressed probability space $(P\mathscr{A}P,\tau/\tau(P))$. Using the Nica-Speicher convolution semi-group, see [MR2266879], Lecture 14, the Voiculescu’s $R$ transform $\mathcal{R}_{PQP}$ of $PQP$ in the compressed space is given by
@@
where $R_{Q}$ is the Voiculescu’s $R$ transform of $Q$ in $(\mathscr{A},\tau)$, (see [MR2266879], page 211). Let
@@
be the shifted Voiculescu R transforms (also known as free cumulant generating functions) of $PQP$ and of $Q$, respectively. Then ([MR2266879], Lecture 16)